# E-values, Multiple Testing and Beyond

### v1

@@ 4.3.2 Simulation studies
We shall compare the finite sample performance of the proposed method with several existing approaches through simulation studies. Throughout, we fix the sample size $n=3,000$ and set the target FDR level at $\alpha_{\mathrm{eBH}}=0.1$. For each experimental setting, we conduct 100 simulations, and report the average FDP (as an estimate of the FDR) and power over the independent simulation runs.
We begin by providing the implementation details of the proposed method. In the GBC procedure, we utilize the rejection rule based on the local FDR under the two group mixture model in Example 3, where we set $f_{0}(p)=\mathbbm{1}\{p\in[0,1]\}$ and $f_{1,i}(p)=(1-\pi_{i})(1-\kappa_{i})p^{-\kappa_{i}}$ with $\kappa_{i}\in(0,1).$ We consider the working models that link $(\pi_{i},\kappa_{i})$ with the external covariates:
where the parameters $\beta_{\pi}=(\beta_{\pi,0},\beta_{\pi,1}^{\top})\in\mathbb{R}^{d+1}$ and $\beta_{\kappa}=(\beta_{\kappa,0},\beta_{\kappa,1}^{\top})\in\mathbb{R}^{d+1}$ can be estimated by maximizing the pseudo-log-likelihood using the EM algorithm. Please refer to Zhang and Chen (2022) for more optimization details. After obtaining the estimates $\hat{\beta}_{\pi}$ and $\hat{\beta}_{\kappa}$ from the EM algorithm, we define
$\hat{\varphi}_{i}(p)=\frac{1}{\hat{\pi}_{i}+(1-\hat{\pi}_{i})(1-\hat{\kappa}_{i})p^{-\hat{\kappa}_{i}}}\leq t,$
where $\hat{\kappa}_{i}=1/\bigl{(}1+\exp(-\beta_{\kappa,0}-\hat{\beta}_{\kappa,1}^{\top}x_{i})\bigr{)}$. Additionally, to speed up the procedure, we propose the following weight that is computationally less expensive:
$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g,j}\}}.$
@@
We fix $\epsilon_{1}=0.1$, $\epsilon_{2}=1\times 10^{-5}$, $G=2$, and $\alpha_{\mathrm{GBC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$. We refer to this method as eBH_GBC for future reference. We compare the proposed method with the following competing methods:
  • BH: The BH procedure (Benjamini and Hochberg, 1995). We implement this method using the p.adjust function in R.
  • IHW_storey: The covariate-powered cross-weighted method with the Storey’s procedure to estimate the null-proportion (Ignatiadis and Huber, 2021). We implement this method using the ihw_bh function in the R package IHWStatsPaper.
  • IHW_betamix: The covariate-powered cross-weighted method with the beta mixture model (Ignatiadis and Huber, 2021). We implement this method using the ihw_betamix_censored function in the R package IHWStatsPaper.
  • AdaPT: The adaptive p-value thresholding procedure (Lei and Fithian, 2018). We implement this method using the adapt_glm function in the R package adaptMT.
  • SABHA: The structure adaptive BH procedure (Li and Barber, 2019). The code was downloaded from the link provided by the original paper.
To illustrate the effect of the covariate, we generate a single covariate $x_{i}$ from the standard normal distribution. Given the value of $x_{i}$, we define $\pi_{i}$ as
$\pi_{i}=\frac{\exp(a_{0}+a_{1}x_{i})}{1+\exp(a_{0}+a_{1}x_{i})},$
@@
where $a_{0}$ and $a_{1}$ determine the baseline signal density and the informativeness of the covariate, respectively. The values of $a_{0}$ and $a_{1}$ are fixed for each simulated dataset. Specifically, we set $a_{0}$ to take on values from the set $\{3.5,1.5,1.5\}$, achieving signal densities of approximately $3\%$, $8\%$, and $18\%$, respectively, representing sparse, medium, and dense signals. Furthermore, we set $a_{1}$ to take on values from the set $\{1.5,2,2.5\}$, representing a less informative, moderately informative, and strongly informative covariate, respectively. The underlying truth $\theta_{i}$ is then simulated based on $\pi_{i}:\theta_{i}\sim\text{Bernoulli}(1-\pi_{i}).$ We next generate the covariate that affects the alternative function $f_{1,i}$. Specifically, we sample another covariate $x_{i}^{\prime}\sim\mathcal{N}(0,1)$ and define
$\eta_{i}=\frac{2\exp(a_{f}x_{i}^{\prime})}{1+\exp(a_{f}x_{i}^{\prime})},$
where we set $a_{f}\in\{0,0.5,1\}$ for no informativeness, less informativeness, and strong informativeness. Then, the z-scores are sampled from
$z_{i}\sim\mathcal{N}(\eta_{i}\mu\theta_{i},1),$
where $\mu$ denotes the signal strength with the values evenly distributed in the interval $[2.5,3.4]$. These $z$-scores are transformed into p-values using the one-sided formula $1-\Phi(z_{i})$. The p-values, along with the corresponding covariates $x_{i}$ and $x_{i}^{\prime}$, serve as the input for the structure adaptive multiple testing methods.
Figure 3: Empirical FDR and power for $a_{f}=1$.
@@
Figure 3 shows the results for $a_{f}=1$. All methods successfully controlled the FDR at the desired level. When the signal is sparse ($a_{0}=3.5$), eBH_GBC is the most powerful method. The SABHA method has the second-best performance, while the two versions of IHW show only slight improvement over the BH procedure. When the covariate is less informative ($a_{1}=1.5$), AdaPT is less powerful than the BH procedure. As the covariate becomes strongly informative ($a_{1}=2.5$), all structure adaptive methods outperform the BH procedure in terms of power. Our proposed method exhibits the highest power in most cases, with only AdaPT surpassing eBH_GBC in power when the signal is dense ($a_{0}=1.5$) and the covariate is strongly informative ($a_{1}=2.5$). The results for the settings with $a_{f}=0$ and $a_{f}=0.5$ are deferred to Appendix C.2.
4.3.3 Real data examples
@@
We analyzed three omics datasets: Airway (Himes et al., 2014), Bottomly (Bottomly et al., 2011), and MWAS (McDonald et al., 2018). The Airway and Bottomly datasets are transcriptomics data obtained from RNA-seq experiments, which have been analyzed previously by Ignatiadis et al. (2016); Lei and Fithian (2018); Zhang and Chen (2022). These two datasets are available in the airway and IHWStatsPaper packages, respectively. For both datasets, we used the logarithm of the “basemean” as the covariate and removed the samples with missing values, leaving us with 18,028 and 11,709 tests, respectively. We obtained the MWAS dataset from the publicly available data of the AmericanGut project (McDonald et al., 2018). We focused on a subset of subjects with ages greater than thirteen and with complete sex and country information. We excluded OTUs (clustered sequencing units representing bacterial species) observed in fewer than ten subjects, resulting in 3,394 OTUs tested using the Wilcoxon rank sum test on normalized abundances. We used the library size of samples as the external covariate.
@@
The results of different methods for target FDR levels ranging from $0.01$ to $0.1$ are presented in Figure 4. AdaPT and eBH_GBC are the two methods that make the most discoveries (except for the MWAS data set with a target FDR level below 0.025). For the airway dataset, AdaPT consistently produces the most discoveries when the target FDR level is high. This is due to the high signal density of this data set. For instance, when the target FDR level is 10%, AdaPT is able to identify 6,053 discoveries out of 18,028 tests. It is worth noting that the proposed method performs similarly to AdaPT when the target FDR level is below 4%. We observe a similar phenomenon for the Bottomly dataset. For the MWAS dataset, AdaPT and eBH_GBC fail to make any discoveries when the target FDR level is 1%. This is a limitation of the BC-based method, which may have reduced power when the signal is very sparse. However, as the target FDR level increases, AdaPT and eBH_GBC quickly surpass the other methods in terms of the number of discoveries. eBH_GBC outperforms AdaPT with more discoveries when the FDR level is above 5%. Overall, eBH_GBC performs comparably to AdaPT.
@@ 5 Discussions
In this paper, we have shown that the BH and BC procedures, as well as their generalized versions, are all equivalent to the e-BH procedure based on some forms of e-values. This simple insight opens up new avenues for constructing multiple testing procedures in different contexts. Specifically, leveraging this idea, one can translate the testing results from different procedures or the same procedure for different subsets of the data into e-values. By aggregating or assembling these e-values, one can obtain a combined set of e-values that contains valuable information about various multiple testing procedures and subsets of the data. We have demonstrated the potential of this strategy in three applications, backed by rigorous theoretical analysis and numerical studies. An important aspect of our proposed methods is the use of data-dependent weights, which are constructed using the leave-one-out technique to aggregate and assemble e-values. Using these weights, we can obtain valid e-values that ensure the corresponding e-BH procedure controls the FDR in finite sample. The e-BH procedure based on the data-dependent weight is often more powerful than the unweighted version, but it comes at the cost of more expensive computation. Through simulations, we have demonstrated that a more cost-friendly version of these weights exists, which achieves nearly identical performance with a much lower computational burden.
@@
Appendix
Appendix A Proof of the main results
We first state the following propositions whose proofs are deferred to Appendix  B. These results will be used frequently in the subsequent proofs of the main theorems.
Proposition 5 (Lemma 6 of Barber et al. [2020]).
@@ Appendix A Proof of the main results
Suppose the assumptions in Proposition 4 hold. Let $T_{i}$ be the threshold for the generalized BC methods when $p_{i}$ is replaced with $\min\{p_{i},1-p_{i}\}$. For any $i$, $j$, if $\max\{\varphi_{i}(1-p_{i}),\varphi_{j}(1-p_{j})\}\leq\max\{T_{i},T_{j}\}$, then we have $T_{i}=T_{j}$.
A.1 Proof of Theorem 1
Proof.
Theorem 1 is a special case of Theorem 2 by choosing $\varphi_{i}$ as the identity function for all $1\leq i\leq n$. ∎
A.2 Proof of Theorem 2
Proof.
Let $\hat{k}$ be the cardinality of $\mathcal{S}_{\mathrm{GBH}}$. If $i\notin\mathcal{S}_{\mathrm{GBH}}$, then $\varphi_{i}(p_{i})>T_{\mathrm{GBH}}$ and thus $e_{i}=0$. This implies that the $i$th hypothesis is not rejected by the e-BH procedure, and hence $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBH}}$. Conversely, if $i\in\mathcal{S}_{\mathrm{GBH}}$, we have $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}\leq\bar{F}^{-1}\left(\frac{\alpha\hat{k}}{n}\right)$, leading to
$e_{i}=\frac{1}{\bar{F}(T_{\mathrm{GBH}})}\geq\frac{n}{\alpha\hat{k}}$
as $\bar{F}$ is strictly increasing. Define $e_{(i)}=n\mathbbm{1}\{q_{(i)}\leq T_{\mathrm{GBH}}\}/\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})$ for $1\leq i\leq n,$ where $q_{(1)}\leq\dots\leq q_{(n)}$ are the order statistics of $\{q_{i}=\varphi_{i}(p_{i})\}^{n}_{i=1}$. Let $\hat{k}$ represent the maximum $i$ for which $q_{(i)}\leq T_{\mathrm{GBH}}$. We get
$e_{(\hat{k})}\geq\frac{n}{\alpha\hat{k}},$ (A.1)
which indicates that $|\mathcal{S}_{\mathrm{eBH}}|\geq\hat{k}$. Because $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBH}}$, it is clear that $\mathcal{S}_{\mathrm{eBH}}=\mathcal{S}_{\mathrm{GBH}}$. ∎
A.3 Proof of Theorem 3
Let $\hat{k}$ be the cardinality of $\mathcal{S}_{\mathrm{GBC}}$. If $i\notin\mathcal{S}_{\mathrm{GBC}}$, then $\varphi_{i}(p_{i})>T_{\mathrm{GBC}}$ and thus $e_{i}=0$. Hence the $i$th hypothesis is not rejected by the e-BH procedure, which implies that $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBC}}$. For the other direction, note that if $i\in\mathcal{S}_{\mathrm{GBC}}$, then $\varphi_{i}(p_{i})\leq T_{\mathrm{GBC}}$ and $1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{\mathrm{GBC}}\}\leq\hat{k}\alpha$. Hence, we have
$e_{i}\geq\frac{n}{\hat{k}\alpha}.$
We sort the e-values in descending order as $e_{(1)}\geq\dots\geq e_{(n)}$. It is clear that $e_{(\hat{k})}\geq n/(\hat{k}\alpha)$. Thus, $|\mathcal{S}_{\mathrm{eBH}}|\geq\hat{k}$, which implies that $\mathcal{S}_{\mathrm{eBH}}=\mathcal{S}_{\mathrm{GBC}}$.
A.4 Proof of Theorem 4
Proof.
Consider the BH procedure and observe that for a given number of rejections $R_{\mathrm{BH}}$, $T_{\mathrm{BH}}=T_{\mathrm{BH}}(R_{\mathrm{BH}})$ is a deterministic function of $R_{\mathrm{BH}}$. Let $R_{\mathrm{BH}}(p_{i}\rightarrow 0)$ be the number of rejections obtained by replacing the p-value $p_{i}$ with 0. Using the above fact and that the weights are independent of the p-values, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}]=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\mathbb{E}\left[\frac{1}{T_{\mathrm{BH}}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\sum_{k=1}^{n}\mathbb{E}\left[\frac{1}{T_{\mathrm{BH}}(k)}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}(k),R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\sum_{k=1}^{n}\mathbb{E}[\mathbbm{1}\{R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}.$
For the BC procedure, denote $\tilde{p}_{i}\coloneqq\min\{p_{i},1-p_{i}\}$ and $\mathbf{p}_{-i}\coloneqq\{p_{1},\dots,p_{i-1},\,\tilde{p}_{i},\,p_{i+1},\dots,p_{n}\}$. By viewing $T_{\mathrm{BC}}$ as a function of the p-values, we define $T_{\mathrm{BC},i}=T_{\mathrm{BC}}(\mathbf{p}_{-i})$. Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. Then, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}\lvert\mathcal{F}_{i}]\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\mathbb{E}[\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}\lvert\mathcal{F}_{i}]\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\right],$
where (i) we have used the fact that $T_{\mathrm{BC}}=T_{\mathrm{BC},i}$ when $p_{i}\leq T_{BC}<0.5$ to get the second equation, (ii) the third equation follows because $T_{\mathrm{BC},i}$ are measurable with respect to $\mathcal{F}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ follows the super-uniform distribution on $[0,1]$, and thus it satisfies Condition (8) under the null.
By Proposition (5), we have
$\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}=\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}.$ (A.2)
If $p_{i}<1-T_{\mathrm{BC},i}$, both sides are equal to 00. If $p_{i}\geq 1-T_{\mathrm{BC},i}$, we claim that $\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}=\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}$. Indeed, if $p_{j}\geq 1-T_{\mathrm{BC},i}$ but $p_{j}<1-T_{\mathrm{BC},j}$, we have $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$. This implies that $\min(p_{i},p_{j})\geq 1-\max\{T_{\mathrm{BC},i},T_{\mathrm{BC},j}\}$. By Proposition 5, we have $T_{\mathrm{BC},i}=T_{\mathrm{BC},j}$, which contradicts with the fact that $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$. The other direction can be proved similarly. Hence we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]\leq n\max_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}\right]\leq n\max_{i}w_{\mathrm{BC},i}.$
By the assumption $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$, we have
$\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i}]\leq\sum_{i=1}^{n}w_{\mathrm{BH},i}+n\max_{j}w_{\mathrm{BC},j}\leq n.$
A.5 Proof of Theorem 5

### v2

@@ 5.2 Simulation Studies
We shall compare the finite sample performance of the proposed method with several existing approaches through simulation studies. Throughout, we fix the sample size $n=3,000$ and set the target FDR level at $\alpha_{\mathrm{eBH}}=0.1$. For each experimental setting, we conduct 100 simulations and report the average FDP (as an estimate of the FDR) and power over the independent simulation runs.
We begin by detailing the implementation of the proposed method. In the FBC procedure, we employ a rejection rule based on the local FDR (Sun and Cai, 2007) within the two-group mixture model framework. Specifically, we propose to use
$\varphi_{i}(p)=\mathrm{Lfdr}_{i}(p)=\frac{\pi_{i}f_{0}(p)}{\pi_{i}f_{0}(p)+(1-\pi_{i})f_{1,i}(p)},$
which represents the posterior probability that the $i$th hypothesis is null given the observed p-value $p$. The literature demonstrates that the rejection rule $\mathbbm{1}\{\varphi_{i}(p_{i})=\mathrm{Lfdr}_{i}(p_{i})\leq t\}$ is optimal in maximizing the expected number of true positives among decision rules that control the marginal FDR at level $\alpha$ (see, e.g., Sun and Cai (2007); Lei and Fithian (2018); Cao et al. (2022)). Additional discussions about local FDR are deferred to Supplement G.1.
Set $f_{0}(p)=\mathbbm{1}\{p\in[0,1]\}$ and $f_{1,i}(p)=(1-\pi_{i})(1-\kappa_{i})p^{-\kappa_{i}}$ with $\kappa_{i}\in(0,1)$, we consider the working models proposed in Zhang and Chen (2022), please refer to Supplement G.1 for more details. After getting $\hat{\pi}_{i}$ and $\hat{\kappa}_{i}$, we define the rejection rule
$\hat{\varphi}_{i}(p)=\frac{\hat{\pi}_{i}}{\hat{\pi}_{i}+(1-\hat{\pi}_{i})(1-\hat{\kappa}_{i})p^{-\hat{\kappa}_{i}}}\leq t.$
We then apply the FBC procedure with the estimated $\hat{\varphi}_{i}$ at the target FDR level $\alpha$. The corresponding e-values are computed via (10), and the weights are obtained from (12). To reduce computational cost, we introduce the following, less expensive weighting scheme:
$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g^{\prime},j}\}}.$
@@
We refer to this method as eBH_FBC for future reference. We compare the proposed method with the following competing methods: BH, IHW_storey, IHW_betamix, AdaPT, and SABHA. The implementation details of these methods are deferred to Supplement G.1.
@@
To illustrate the effect of the covariate, we generate a single covariate $x_{i}$ from the standard normal distribution. Given the value of $x_{i}$, we define $\pi_{i}$ as $\pi_{i}=\exp(a_{0}+a_{1}x_{i})/\bigl{(}1+\exp(a_{0}+a_{1}x_{i})\bigr{)}$, where $a_{0}$ and $a_{1}$ determine the baseline signal density and the informativeness of the covariate, respectively. The values of $a_{0}$ and $a_{1}$ are fixed for each simulated dataset. Specifically, we set $a_{0}$ to take on values from the set $\{3.5,2.5,1.5\}$, achieving signal densities of approximately $3\%$, $8\%$, and $18\%$, respectively, representing sparse, medium, and dense signals. Furthermore, we set $a_{1}$ to take on values from the set $\{1.5,2,2.5\}$, representing a less informative, moderately informative, and strongly informative covariate, respectively. The underlying truth $\theta_{i}$ is then simulated based on $\pi_{i}:\theta_{i}\sim\text{Bernoulli}(1-\pi_{i}).$ We next generate the covariate that affects the alternative function $f_{1,i}$. Specifically, we sample another covariate $x_{i}^{\prime}\sim\mathcal{N}(0,1)$ and define $\eta_{i}=2\exp(a_{f}x_{i}^{\prime})/\bigl{(}1+\exp(a_{f}x_{i}^{\prime})\bigr{)}$, where we set $a_{f}\in\{0,0.5,1\}$ for no informativeness, less informativeness, and strong informativeness. Then, the z-scores are sampled from $z_{i}\sim\mathcal{N}(\eta_{i}\mu\theta_{i},1)$, where $\mu$ denotes the signal strength with the values evenly distributed in the interval $[2.5,3.4]$. These $z$-scores are transformed into p-values using the one-sided formula $1-\Phi(z_{i})$. The p-values, along with the corresponding covariates $x_{i}$ and $x_{i}^{\prime}$, serve as the input for the structure-adaptive multiple testing methods.
Figure 3: Empirical FDR and power with $a_{f}=1$. Signal sparsity is controlled by setting $a_{0}\in\{3.5,2.5,1.5\}$, giving rise to sparse, moderate, and dense alternatives, respectively. Covariate informativeness is tuned via $a_{1}\in\{1.5,2,2.5\}$, corresponding to weak, moderate, and strong auxiliary signals.
@@
Figure 3 presents the results for $a_{f}=1$. All methods successfully controlled the FDR at the desired level. When the signal is sparse ($a_{0}=3.5$), eBH_FBC is the most powerful method. The SABHA method exhibits the second-best performance, while both versions of IHW show only slight improvements over the BH procedure. When the covariate is less informative ($a_{1}=1.5$), AdaPT is less powerful than the BH procedure. As the covariate becomes strongly informative ($a_{1}=2.5$), all structure-adaptive methods outperform the BH procedure in terms of power. Our proposed method demonstrates the highest power in most cases, with AdaPT surpassing eBH_FBC in power when the signal is dense ($a_{0}=1.5$) and the covariate is strongly informative ($a_{1}=2.5$). The results for settings with $a_{f}=0$ and $a_{f}=0.5$ are deferred to Supplement G.2.
To demonstrate the stability of each method in structure adaptive multiple testing, particularly the consistency of results when the data are randomly generated from the same distribution, we plot the variance of the FDP in Figure G.1 in Supplement G.2. The figure shows that as the signal becomes dense or the covariate contains more information, the variance of all methods decreases. Notably, the proposed method exhibits a smaller variance compared to AdaPT. Additionally, Table G.1 in Supplement G.2 displays the running time for each method to compare their computational efficiency. We focus on a simulation setting with $a_{0}=1.5$, $a_{1}=2$, $a_{f}=1$, and $\mu=3$. We conducted 100 simulation runs for each method and reported the average time taken to complete the analysis. The results indicate that our method is approximately ten times faster than AdaPT.
5.2.1 Real-Data Examples
@@
We analyzed three omics datasets: Airway (Himes et al., 2014), Bottomly (Bottomly et al., 2011), and MWAS (McDonald et al., 2018). The Airway and Bottomly datasets are transcriptomics data obtained from RNA-seq experiments. For both datasets, we used the logarithm of the ‘basemean” as the covariate and removed the samples with missing values, leaving us with 18,028 and 11,709 tests, respectively. We obtained the MWAS dataset from the publicly available data of the American Gut project (McDonald et al., 2018). We focused on a subset of subjects with ages greater than thirteen and with complete sex and country information. We excluded OTUs observed in fewer than ten subjects, resulting in 3,394 OTUs tested using the Wilcoxon rank sum test on normalized abundances. We used the library size of samples as the external covariate.
The results of different methods for target FDR levels ranging from $0.01$ to $0.1$ are presented in Figure 4. AdaPT and eBH_FBC are the two methods that make the most discoveries (except for the MWAS data set with a target FDR level below 0.025). For the airway dataset, AdaPT consistently produces the most discoveries at higher target FDR levels, which is due to the high signal density of this dataset. For instance, when the target FDR level is 10%, AdaPT is able to identify 6,053 discoveries out of 18,028 tests. It is worth noting that the proposed method performs similarly to AdaPT when the target FDR level is below 4%. We observe a similar phenomenon for the Bottomly dataset. For the MWAS dataset, both AdaPT and eBH_FBC fail to make any discoveries when the target FDR level is set to 1%. This is a limitation of the BC-type method, which may have reduced power when the signal is very sparse. However, as the target FDR level increases, AdaPT and eBH_FBC quickly outperform the other methods in terms of the number of discoveries. eBH_FBC outperforms AdaPT with more discoveries when the FDR level is above 5%. Overall, eBH_FBC performs comparably to AdaPT.
@@ 6 Discussions
Motivated by the recent findings in Li and Zhang (2025), we transform testing results from different procedures or data subsets into e-values. By aggregating these e-values, we obtain a combined set of e-values that captures information across procedures or partitions. A key feature of our method is the use of data-dependent weights, constructed via a leave-one-out approach, to ensure the resulting e-values yield finite-sample FDR control under the e-BH procedure. This weighted version is often more powerful than its unweighted counterpart. Simulations further reveal that a computationally efficient approximation of the weights achieves comparable performance.
@@
SUPPLEMENTARY MATERIAL
Supplement:
Including all the proofs, additional discussions, and numerical results.
Appendices
Appendix A Preliminaries
A.1 Unified Form of Multiple Testing Procedures
Table A.1 summarizes the specifications of $m(t)$ and $R_{i}(t)$ for different multiple testing procedures.
Method $m(t)$ $R_{i}(t)$
BH ntnt $\mathbbm{1}\{p_{i}\leq t\}$
ST $n\pi_{0}^{\lambda}t$ $\mathbbm{1}\{p_{i}\leq t\}$
BC $1+\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\geq 1-t\}$ $\mathbbm{1}\{p_{i}\leq t\}$
FBH $ng(t)$ $\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$
FBC $1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}$ $\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$
Table A.1: Definitions of $m(t)$ and $R_{i}(t)$ for various multiple testing procedures.
Appendix B Proofs of Main Results
We begin by stating the following propositions, whose proofs are deferred to Appendix C to ensure self-containment. These results will be frequently utilized in the subsequent proofs of the main theorems.
Proposition B.1 (Lemma 6 of Barber et al. [2020]).
@@ Appendix B Proofs of Main Results
Suppose that the null p-values are mutually independent, are independent of the alternative p-values, and satisfy Condition (6). Let $T_{i}$ be the threshold for the generalized BC methods when $p_{i}$ is replaced with $\min\{p_{i},1-p_{i}\}$. For any $i$, $j$, if $\max\{\varphi_{i}(1-p_{i}),\varphi_{j}(1-p_{j})\}\leq\max\{T_{i},T_{j}\}$, then we have $T_{i}=T_{j}$.
B.1 Proof of Theorem 2